# WiLoR: End-to-end 3D Hand Localization and Reconstruction in-the-wild

### v1

@@ 5 Experiments
In this section, we first evaluate the proposed hand detection network using established benchmarks to assess its performance. Subsequently, we conduct an extensive qualitative and quantitative analysis of the proposed 3D hand pose estimation method. Finally, we demonstrate the critical role of precise hand localization in enhancing the accuracy of 4D hand reconstruction.
@@ 5.1 Evaluation of Hand Detection and Localization
Training. We train the proposed hand detection network using the curated WHIM dataset that consists of over 2M in-the-wild images of multiple hands and scales. To further boost the generalization and robustness of our network, we follow several data augmentations during training. Particularly, we introduce random rotations in the range of $[-60^{\circ},60^{\circ}]$ and translations in the range of $[-0.1,0.1]$ along with random masking and cropping in the image. Additionally, in each training batch, we follow mosaic and mixup augmentation, which significantly affects the robustness to diverse hand scales.
@@
In Tab. 1, we evaluate the proposed and the baseline methods on three datasets: the proposed WHIM dataset along with the benchmark Coco-WholeBody [32] and Oxford-Hands [48] dataset. All experiments were conducted on a NVIDIA RTX 4090 GPU. As can be easily seen, the proposed detector can run in over than 130 FPS while achieving up to 21% improvement on AP compared to previous state-of-the-art models. In addition, compared to previous state-of-the-art model, ContactHands [52], the proposed detector is 45×45$\times$45 × faster and has 32×32$\times$32 × reduced model size which enables the utilization of the proposed detector in mobile applications and heavy pipelines without posing any significant overhead. It is important to note that despite the varying resolutions and hand scales across the three datasets, the proposed model consistently outperforms the baseline methods. This is particularly evident on the COCO-WholeBody dataset [32], an extension of the COCO dataset that includes full-body images, where the hands are relatively small compared to the overall image size.
@@
Ablation. The efficiency and accuracy of the proposed hand detection method are mainly attributed to the selection of the backbone architecture and the utilization of a large-scale training dataset. We further evaluate the contribution of each component using an ablation study on OxfordHands [48] and WHIM datasets where we utilized different backbone networks and training datasets. As can be seen from Tab. 2, trained the detection network with different backbones, apart from achieving similar detection performance, significantly degraded the inference speed of the network. The importance of the proposed large-scale in-the-wild WHIM dataset is also validated in Tab. 2, where we can observe a significant performance drop when the model was trained with significantly less data, e.g., Proposed w. 0.25M, Proposed w. 0.5M, Proposed w. 1M. An interesting observation highlighting the versatility of the WHIM dataset is that the proposed model achieves better performance when trained on WHIM compared to the OxfordHands dataset. Finally, we evaluate the contribution of the proposed augmentation strategy and the use of landmark regression loss. The augmentation strategy significantly contributes to cross-dataset generalization, achieving 14% increase on mAP. Similarly, we can observe that incorporating landmark regression loss enhances the detector’s precision, leading to more robust detections.
@@
Method Size (Mb) FPS AP0.5 mAP AP0.5 mAP
@@
5.2 3D Hand Pose Estimation
Training. Following [6, 40, 61], we trained the proposed hand regressor using a combination of datasets to improve robustness to diverse poses, illuminations and occlusions. Particularly, we utilized a set of datasets containing both 2D and 3D annotations namely FreiHAND [100], HO3D [24], MTC [89], RHD [99], InterHand2.6M [50], H2O3D [24], DEX YCB [8], COCO WholeBody [32], Halpe [17] MPII NZSL [75], BEDLAM [4], ARCTIC [16], Re:InterHand  [51] and Hot3D  [3]. In total we utilized 4.2M images, 55% more than previous state-of-the-art. Including egocentric datasets such as ARCTIC [16] and Hot3D [3] significantly improve the reconstruction performance of the proposed method on egocentric views.
Evaluation. We compare the proposed method with state-of-the-art methods including METRO [41], Mesh Graphormer [42], AMVUR [31], MobRecon [10], HaMeR [61] and SimpleHand [96]. In Tab. 3 and Tab. 4 we report the reconstruction results the popular benchmark FreiHAND [100] and HO3Dv2 [24] datasets. Following the common protocol [100], we measure the reconstruction performance in terms of Procrustes Aligned Mean per Joint and Vertex Error (PA-MPJPE, PA-MPVPE) along with the fraction of poses with less than 5mm and 15mm error (F@5, F@15). Additionally, we report Area Under the Curve for 3D joints and vertices ($\textrm{AUC}_{\textrm{J}}$, $\textrm{AUC}_{\textrm{V}}$) for HO3D dataset. As can be observed the proposed method achieves state-of-the-art performance and outperforms previous methods under all metrics on both benchmark datasets, which can be further validated in Fig. 6.
@@
5.3 Dynamic Reconstruction
A key challenge for 3D pose estimation methods is to achieve stable and robust 4D reconstructions without being trained on a dynamic setting [74]. Traditionally, methods for 3D pose estimation from single image suffer from low temporal coherence and jittering effects across frames and can not generalize well to videos, setting a huge burden on their generalization to real-world reconstruction. To effectively evaluate the temporal coherence of the proposed method, we reconstruct frame-wise the 4D sequence and measure the jittering between frames. In particular, we calculate the mean per frame Euclidean distance of the 3D vertices (MPFVE) and joints (MPFJE) between consecutive frames. Additionally, similar to [74], we measure the jerk (Jitter) of the 3D hand joints motion along with the global Root Translation Error (RTE) that measures the displacement of the wrist across frames. In Tab. 6, we report the reconstruction results fro the best performing methods on HO3D [24] dataset. WiLoR outperforms baseline methods in temporal coherence without relying on any temporal module based on the robust stability of the detections.
@@ 5.3 Dynamic Reconstruction
Table 6: Reconstruction of dynamic 3D Hands. We evaluate the temporal coherence and the jittering of the reconstruction for the proposed and the baseline methods on the HO3D dataset. We include for reference the ground truth values.
@@ 6 Conclusion
In this work, we propose the first full-stack hand detection and 3D pose estimation framework. Using a large-scale in-the-wild dataset we train a light-weight yet highly accurate hand detector model that can robustly detect hands under different occlusions and illuminations at over 110 FPS. Additionally, we propose a high fidelity 3D hand pose estimation model build on top of our novel refinement module, that overcomes the limitations of previous methods and mitigates the alignment issues of previous methods. Under a series of experiments, we showcase that the proposed method can outperform previous state-of-the-art on two benchmark datasets and show robust performance on challenging cases.
@@
Acknowledgements S. Zafeiriou was supported by EPSRC Project DEFORM (EP/S010203/1) and GNOMON (EP/X011364). R.A. Potamias was supported by EPSRC Project GNOMON (EP/X011364).

### v2

@@ 5 Experiments
In this section we first evaluate the proposed hand detection network using established benchmarks to assess its performance. Next, we conduct an extensive qualitative and quantitative analysis of the proposed 3D hand pose estimation method. Finally, we demonstrate the critical role of precise hand localization in the accuracy of 4D hand reconstruction.
@@ 5.1 Evaluation of Hand Detection and Localization
Training. We train the proposed hand detection network using the curated WHIM dataset that consists of over 2M in-the-wild images of multiple hands and scales. To further boost the generalization and robustness of our network, we follow several data augmentations during training. Particularly, we introduce random rotations in the range of $[-60^{\circ},60^{\circ}]$ and translations in the range of $[-0.1,0.1]$ along with random masking and cropping of the image. Additionally, in each training batch, we follow mosaic and mixup augmentation, which significantly affects the robustness to diverse hand scales.
@@
In Tab. 1, we evaluate the proposed and the baseline methods on three datasets: the proposed WHIM dataset along with the benchmark Coco-WholeBody [33] and Oxford-Hands [50] dataset. All experiments were conducted on a NVIDIA RTX 4090 GPU. As can be easily seen, the proposed medium size detector, Proposed-M, can run at more than 130 FPS while the small version, Proposed-S, can achieve up to 175 FPS, improving the mAP metric, on average, by 26% compared to previous state-of-the-art model. In addition, compared to previous state-of-the-art model, ContactHands [54], the proposed detector is 45×45$\times$45 × faster and has 32×32$\times$32 × reduced model size which enables the utilization of the proposed detector in mobile applications and heavy pipelines without posing any significant overhead. It is important to note that despite the varying resolutions and hand scales across the three datasets, the proposed model consistently outperforms the baseline methods. This is particularly evident on the COCO-WholeBody dataset [33], an extension of the COCO dataset that includes full-body images, where the hands are relatively small compared to the overall image size.
@@
Ablation. The efficiency and accuracy of the proposed hand detection method are mainly attributed to the selection of the backbone architecture and the utilization of a large-scale training dataset. We further evaluate the contribution of each component using an ablation study on OxfordHands [50] and WHIM datasets where we utilized different backbone networks and training datasets. As can be observed from Tab. 2, training the detection network with different backbones, apart from achieving similar detection performance, significantly degrades the inference speed of the network. The importance of the proposed large-scale in-the-wild WHIM dataset is also validated in Tab. 2, where we can observe a significant performance drop when the model was trained with significantly less data, e.g., Proposed w. 0.25M, Proposed w. 0.5M, Proposed w. 1M. An interesting observation highlighting the versatility of the WHIM dataset is that the proposed model achieves better performance when trained on WHIM compared to the OxfordHands dataset. Finally, we evaluate the contribution of the proposed augmentation strategy and the use of landmark regression loss. The augmentation strategy significantly contributes to cross-dataset generalization, achieving 14% increase on mAP. Similarly, we can observe that incorporating landmark regression loss enhances the detector’s precision, leading to more robust detections.
@@
Method Size (Mb) $\downarrow$ FPS $\uparrow$ AP0.5 $\uparrow$ mAP $\uparrow$ AP0.5 $\uparrow$ mAP $\uparrow$
@@
5.2 Evaluation of 3D Hand Pose Estimation
Training. Following [7, 42, 63], we trained the proposed hand regressor using a combination of datasets to improve robustness to diverse poses, illuminations and occlusions. Particularly, we utilized a set of datasets containing both 2D and 3D annotations namely FreiHAND [107], HO3D [25], MTC [92], RHD [106], InterHand2.6M [52], H2O3D [25], DEX YCB [9], COCO WholeBody [33], Halpe [18] MPII NZSL [78], BEDLAM [4], ARCTIC [17], Re:InterHand  [53] and Hot3D  [3]. In total we utilized 4.2M images, 55% more than previous state-of-the-art.
Evaluation. To compare the proposed method we employ with state-of-the-art methods including METRO [43], Mesh Graphormer [44], AMVUR [32], MobRecon [11], HaMeR [63] and SimpleHand [103]. In Tab. 3 and Tab. 4 we report the reconstruction results the popular benchmark FreiHAND [107] and HO3Dv2 [25] datasets. Following the common protocol [107], we measure the reconstruction performance in terms of Procrustes Aligned Mean per Joint and Vertex Error (PA-MPJPE, PA-MPVPE) along with the fraction of poses with less than 5mm and 15mm error (F@5, F@15). Additionally, we report Area Under the Curve for 3D joints and vertices ($\textrm{AUC}_{\textrm{J}}$, $\textrm{AUC}_{\textrm{V}}$) for HO3D dataset. The proposed method achieves state-of-the-art performance and outperforms previous methods under all metrics on both benchmark datasets, which can be further validated qualitatively in Fig. 6. Leveraging the image-aligned features of the refinement module, WiLoR achieves high fidelity reconstructions even in challenging articulations.
@@ 5.2 Evaluation of 3D Hand Pose Estimation
5.3 Evaluation of Dynamic Reconstruction
A key challenge for 3D pose estimation methods is to achieve stable and robust 4D reconstructions without being trained using a dynamic setting [77]. Traditionally, methods for 3D pose estimation from single image suffer from low temporal coherence and jittering effects across frames, setting a huge burden on their generalization to real-world video reconstruction. To effectively evaluate the temporal coherence of the proposed method, we reconstructed frame-wise a 4D sequence and measure the jittering between frames. In particular, we calculate the mean per frame Euclidean distance of the 3D vertices (MPFVE) and joints (MPFJE) between consecutive frames. Additionally, similar to [77], we measure the jerk (Jitter) of the 3D hand joints motion along with the global Root Translation Error (RTE) that measures the displacement of the wrist across frames. In Tab. 6, we report the reconstruction results for the best performing methods on HO3D [25] dataset. WiLoR outperforms baseline methods in temporal coherence without relying on any temporal module based on the robust stability of the detections. We refer the reader to the supplementary material for qualitative video results.
@@ 5.3 Evaluation of Dynamic Reconstruction
Table 6: Reconstruction of dynamic 3D Hands. We evaluate the temporal coherence and the jittering of the reconstruction for the proposed and the baseline methods on the HO3D dataset.
@@ 6 Conclusion
In this work, we present WiLoR, the first full-stack hand detection and 3D pose estimation framework. Using a large-scale in-the-wild dataset we train a light-weight yet highly accurate hand detector model that can robustly detect hands under different occlusions and illuminations at over 130 FPS. Additionally, we propose a high fidelity 3D hand pose estimation model built on top of our novel refinement module, that overcomes the limitations of previous methods and mitigates the alignment issues of previous methods. Under a series of experiments, we showcase that WiLoR outperforms previous state-of-the-art methods on two benchmark datasets and show robust performance on challenging cases. WiLoR establishes a comprehensive solution for multi-hand detection, localization and 3D reconstruction.
@@
Acknowledgements S. Zafeiriou was supported by Turing AI Fellowship (EP/Z534699/1) and GNOMON (EP/X011364). R.A. Potamias was supported by EPSRC Project GNOMON (EP/X011364).
Supplementary Material
7 Implementation Details
In this section we report the training details of the hand detection and the hand pose estimation models.
7.1 Hand Detection and Localization
To train the detector model we use WHIM dataset that comprises of over 2M in the wild images from daily activities. We train WiLoR detector with Adam optimizer for 200 epochs with early stopping if there is no loss decrease for over 30 epochs. Initiate the training with a learning rate of 0.01 and linearly decrease to $61e-$ 6 for the last 30 epochs of the training. We trained the model for three weeks using two NVIDIA RTX 4090 and a batch size of 256. To weight the different losses we set $\lambda_{0}=0.5$ for the classification loss, $\lambda_{1}=1.5$ for the distribution focal length loss, $\lambda_{2}=15$ for the bounding box loss, $\lambda_{3}=10$ for the keypoints loss. We use random mosaic augmentations with probability 0.7, random rotations between $[-60^{o},60^{o}]$ and random image scaling between [0.5, 1].
7.2 Hand Pose Estimation
We build our hand pose estimation method on top of a ViT-Large backbone with pre-trained weights from ViTPose [94], with a hidden dimension of 1280. Apart from the image patches, we use three additional learnable tokens that correspond to hand pose, shape and camera translation and scale. We initialize the tokens with the mean pose, shape and camera parameters from the training set. Using a set of fully-connected layers, we map the output tokens to 96 MANO pose parameters (15 joint rotations + 1 global orientation represented in 6d rotation format  [102]), 10 MANO shape parameters and a 3D camera translation. We then reshape the output image tokens to a $16\times 12$ image form, and perform two sets of upsamplings using deconvolutions. At each upsampling step we reduce the feature by 2 times. Using the initially estimated camera parameters we project the rough MANO estimation to the feature maps and sample a set of multi-scale per-vertex image-aligned features. The concatenated set of features is then aggregated and regressed from a set of fully-connected layers that predict the pose, shape and camera residuals. We train the model for 1000 epochs using Adam optimizer with an initial learning rate of 1e-5 and a weight decay of 1e-4. Similar to the hand detector, we apply random scaling, rotations and color jitter during training. Similar to  [63], to balance the losses we set $\lambda_{3D}=0.05$, $\lambda_{2D}=0.01$, $\lambda_{pose}=0.001$, $\lambda_{shape}=0.0005$ and $\lambda_{adv}=0.0005$.
8 Comparison with existing datasets
In contrast to 3D hand pose estimation methods that utilizes images of tightly cropped hands, to train a powerful hand detector network, it is required to create a dataset that contains images with multiple hands under different occlusions, views, illuminations and skin tones. Bellow we compare WHIM with such available datasets. WHIM is 100$\times$ larger than previous in-the-wild multi-hand datasets.
Dataset #Img Annotations Egocentric Third-Person Objects Real 3D
OxfordHands 13K Manual
MPI-HP 25K Manual
Coco-Whole 200K Manual
BEDLAM 380K GT
AGORA 18K GT
ContactHands 21K Manual
CocoHands 25K Manual
BodyHands 20K Manual
WHIM 2M Auto
Table 7: Comparison with existing hand datasets. WHIM is 100$\times$ larger than previous multi-hand dataset.s
9 Limitations
Although achieving state-of-the-art performance on both 3D hand pose estimation and hand detection tasks, WiLoR still fails to recover challenging cases. Despite being trained on a large-scale dataset, the data distribution is still limited to ‘common’ hand poses and appearances, failing to generalize to samples far from the trained distribution. As can be seen in Fig. 7 WiLoR can fail under extreme finger poses and can also fail to detect hands in crowded environments. Creating a synthetic dataset with diverse hand poses and photorealistic hands could help mitigate these issues  [67]. Additionally, since WiLoR employs a bottom-up reconstruction strategy, interactions and contacts between hands may not be adequately captured in 3D space. In scenarios where accurate hand contact estimation is crucial  [2, 108], incorporating additional interaction constraints [93] may be necessary. Finally, WiLoR estimates 3D hand poses in camera space, which may lead to inaccurate assumptions about the overall 3D scene. Adapting WiLoR with a 3D metric foundational model [96, 99] could enable more accurate 3D reconstruction in world space.
Figure 7: Failure Cases. WiLoR can still fail to reconstruct complex finger poses or detect small hands in crowded environment.
10 Training Datasets
To train our hand pose estimation module we use a combination of datasets to enforce the generalization of the model. In particular, we use 14 datasets with both 2D and 3D annotations, from three major categories: controlled environment hand images, hand-object interaction, in-the-wild and synthetic datasets, resulting in 4.2M images total:
  • FreiHAND [107] is a common 3D hand pose estimation dataset composed of 132K images of indoor, outdoor, and synthetic scenes. It provides both 3D hand and 2D keypoint annotations.
  • MTC [92] is a subset of Panoptic Studio Dataset [34] that contains 360K multi-view images in a studio environment. The dataset provides both 3D hand and 2D keypoint annotations.
  • InterHand2.6M [52] is a large scale environment from light stage environment that contains hand articulations from 27 different subjects and 80 different cameras. The dataset provides both 3D hand and 2D keypoint annotations.
To increase the generalization of WiLoR under severe occlusions we include several datasets where hands interact with objects.
  • HO3D [25] provides a hand-object dataset with over than 120K images from multi-view cameras of hands interacting with objects. It is used as one of the main benchmarks for hand and object reconstruction. Images were captured in an lab environment setting. It provides both 3D hand and 2D keypoint annotations.
  • H2O3D [25] contains over 60K images from five multi-view cameras of hands interacting with objects. In contrast to HO3D, each subject is interacting with the object using two hands which increases the occlusions of the hands. It provides both 3D hand and 2D keypoint annotations.
  • DEX YCB [9] similar to HO3D, DEX YCB is a benchmark dataset for hand object reconstruction. It contains over than 500K multi-view images from 10 objects grasping objects. The dataset provides both 3D hand and 2D keypoint annotations.
  • ARCTIC [17] is a large scale dataset of bimanual hand-object manipulations containing over than 400K images from both egocentric and third-person views. It contains both single and dual hand manipulations along with accurate 3D and 2D hand annotations.
  • Hot3D  [3] is a recent egocentric dataset from daily activities that includes a high degree of occlusions and can significantly enhance the performance of WiLoR in egocentric scenarios. It contains both 3D hand and 2D keypoint annotations.
Additionally, we include four synthetic datasets that provide accurate ground truth 2D and 3D annotations:
  • RHD [106] is amongst the first synthetic datasets of hands rendered under different illumination patterns. The dataset is composed of 62K images with accurate 3D and 2D annotations.
  • Re:InterHand  [53] is synthetic dataset that extends InterHand2.6M by rendering hands under different illuminations and environments to bridge the gap between studio setup and in-the-wild images. The dataset provides both 3D hand and 2D keypoint annotations.
  • BEDLAM [4] is a large scale full body synthetic dataset, that has proven extremely effective in whole body reconstruction tasks [7]. We use BEDLAM and randomly crop regions around the human hands to augment the training data. We use over 500K image crops. The dataset provides both 3D hand and 2D keypoint annotations.
Finally, we include in-the-wild datasets that contain only 2D information, but can effectively boost the generalization performance of WiLoR in in the while scenarios.
  • COCO WholeBody [33] is a subset of COCO dataset and one of the main benchmarks for body pose estimation. It contains over than 100K in the wild images with humans participating in different activities. It provides 2D hand keypoint annotations.
  • Halpe [18] is a full body in-the-wild dataset, composed of over than 40K images will 2D keypoint annotations. The dataset contains 21 keypoints for each hand.
  • MPII NZSL [78] is a common benchmark for human body pose estimation, containing a set of in-the-wild, synthetic and lab environment images. The dataset contain 15K images with 2D keypoint hand annotations.
11 Temporal Coherence
In the project page we provide several videos that demonstrate the temporal coherence of WiLoR in challenging scenarios such as kneading dough or playing guitar. Despite being trained on single images, WiLoR can provide smooth reconstructions given its stable and robust detections.